Problem 5 If z and y are positive, then > V2+y Answer:
Answer #1

So, for verifying the equation, it is given that x and y are positive integers,

So let us assume x = 8 and y = 6,

therefore, x + y = 8 + 6 = 14

on the RHS,

√(x2 + y2)

= √(64+36)

= √100

= 10 < x + y proved.

Please find the screenshot of the proof in python programming below:

С https://www.onlinegdb.com/online-python-compiler Run DebugStop C Share SaveBeautify Language Python 3 OnlineGDB beta main.